OpenStudy (anonymous):

Jason and Kyle both choose a number from 1 to 10 at random. What is the probability that both numbers are odd?

Parth (parthkohli):

Probability(Odd number){1,3,5,7,9} = 5/10 = 1/2 Since we are having a word 'and','both' or 'simultaneous', we will MULTIPLY the probabilities. \(\Large \color{MidnightBlue}{\Rightarrow {1 \over 2} * {1 \over 2} }\)

OpenStudy (anonymous):

verry insightful thanks